# Homological Filling and Minimal Varifolds in Four-Dimensional Einstein Manifolds

### v1

@@ 2.4. Bubble–tree decomposition
There exist integers $N_{E}(v,D)$, $k_{E}(v,D)$ and a decomposition
@@
    (Bodies) If $x\in\mathcal{B}^{i}_{j}$ then $r_{h}(x)\geq r_{0}(v,D)\,\operatorname{diam}(\mathcal{B}^{i}_{j}),$ where $r_{h}$ is the harmonic radius and $r_{0}(v,D)$ is as in Corollary 2.6.
@@
    (Necks) Each $\mathcal{N}^{i}_{j}$ is diffeomorphic to an annulus in $\mathbb{R}^{4}/\Gamma^{i}_{j}$, with $|\Gamma^{i}_{j}|\leq C_{\Gamma}(\mathbf{A}(v,D))$ for some explicit function $C_{\Gamma}$.
@@
This a specialization of Cheeger–Naber’s bubble–tree construction [8, Theorem 8.64] to the Einstein class, using the $\varepsilon$-regularity Theorem 2.5 and the global $L^{2}$-bound of Theorem 2.3. We denote the adjacency relations between bodies and necks as a rooted tree with root $\mathcal{B}^{1}$ and denote by $\mathcal{T}^{i}_{j}$ the subtree with root $\mathcal{B}^{i}_{j}$.
@@ 3.1. Homological filling functions and stationary varifolds
Thus, in order to bound $A_{\min}(M)$ from above, it suffices to obtain a linear upper bound for $\operatorname{FH}_{1}(l)$ in terms of the geometric data of $(M,g)$. The rest of this section is devoted to establishing such a bound in the Einstein setting.
@@ 3.2. A homological filling inequality for Einstein manifolds
A general $1$-cycle $C$ in $M$ will first be homologically replaced by a sum of cycles supported in single bodies, then approximated by a cycle in $\Gamma$, then pushed to a simplicial cycle in $\mathfrak{N}$ and filled combinatorially using Lemma 3.7. Each $2$-simplex in $\mathfrak{N}$ corresponds to a geodesic triangle in $\Gamma$, which can be filled by a $2$-chain of uniformly bounded area. This yields the desired linear filling inequality.
@@
where $r_{h}$ is the harmonic radius and $r_{0}(v,D)$ is the scale from Corollary 2.6. Set
@@
The number of sets in $\mathcal{O}$ is bounded by $|\mathcal{O}|\leq\tilde{N}_{E}(v,D)$ for some explicit $\tilde{N}_{E}(v,D)$ depending only on the bubble–tree complexity $N_{E}(v,D)$, $k_{E}(v,D)$ and the covering constants $N_{1}(v,D)$, $N_{2}(v,D)$; see also [23, Sec. 4] for the general Ricci-bounded case.
@@
for some explicit constant $C(v,D)$ depending only on the geometric data. This construction is identical to that in [23, Sec. 4], with the only difference that the scales $R^{i}_{j}$ and $\rho$ are now tied to the Einstein harmonic-radius scale $r_{0}(v,D)$.
@@ 3.3. Local fillings
where $h_{E}(v,D)$ depends only on an upper bound for $|\Gamma|$ (hence on $\mathbf{A}(v,D)$ and $C_{S}(v,D)$).
@@
The argument is exactly as in [23, Sec. 4.2]. Via $\Phi^{-1}$ we view $C$ as a $1$-cycle in the annulus $A_{\rho,2\rho}(0)\subset\mathbb{R}^{4}/\Gamma$. Radial projection onto a sphere $S_{\lambda}/\Gamma$ with $\lambda\in(\rho,2\rho)$ sweeps out a $2$-chain of area $\lesssim\rho\,\operatorname{mass}_{1}(C)$. The projected cycle is a $1$-cycle in $S^{3}/\Gamma$, whose first homology is finite with generators represented by loops of length $\leq 2\pi\lambda$. Replacing the projected cycle by a representative of its homology class supported in a bounded number of such generators yields $C^{\prime}$ with $\operatorname{mass}_{1}(C^{\prime})\leq h_{E}(v,D)\,\lambda$. Pushing forward by $\Phi$ and controlling the distortion inside the annulus completes the proof. ∎
@@ 3.4. Approximation by $\Gamma$ and improved combinatorial filling
The proof is identical to [23, Lemma 4.5]. One partitions $C$ into short arcs contained in the balls of the cover $\mathcal{O}(\mathcal{B})$, with length $\leq R^{i}_{j}$. Inside each ball one joins the endpoints of an arc to the corresponding center along edges of $\Gamma$. The difference between the original arc and the resulting path in $\Gamma$ lies in a harmonic ball and hence can be filled using Lemma 3.3. Summing over all arcs gives the desired estimates. ∎
The idea of controlling the combinatorial filling via integer solutions of linear systems was introduced in [19]. We are going to show that it is possible to obtain an upper bound of the number of simplicies needed to fill a simplicial cycle which depends linearly on the size of the cycle. Note that this is independent of the dimension of the simplicial complex.
@@
As in [23, Lemma 4.6]. If all three vertices lie inside a single body, the triangle lies in a harmonic ball $B_{r_{h}/2}$ and can be filled using Lemma 3.3. If they lie inside a single neck, the triangle lies in the union of at most two trapezoids contained in $\bar{T}_{i,j,\ell}$, and can be filled by a cone in the annulus, with area $\lesssim\rho^{2}$. Mixed configurations are reduced to these two cases. All constants can be controlled in terms of $r_{0}(v,D)$ and $D$. ∎
@@ 3.5. Decomposition and the Einstein filling inequality
As in [23, Sec. 4.3], one cuts $C$ whenever it crosses a boundary component $\partial\mathcal{B}\cap\mathcal{N}$ and reassembles the resulting arcs into cycles supported in individual bodies or necks. The number of such regions is bounded by the bubble–tree complexity $N_{E}(v,D)\,k_{E}(v,D)$, and the total length increases by at most a uniform multiplicative factor $C_{0}(v,D)$, coming from bounded overlap of the decomposition and the uniform control on the diameter of bodies and necks. ∎
@@
$\partial E_{0}=C-C^{\mathrm{body}},\ \operatorname{mass}_{2}(E_{0})\leq C_{2}(v,D)\,\operatorname{mass}_{1}(C)+C_{3}(v,D),$
@@
The cycle $\widehat{C}$ maps under $f$ to a simplicial $1$-cycle f(C^)f($\widehat${C}) in the nerve $\mathfrak{N}$. Let $n_{0}$ be the number of vertices of $\mathfrak{N}$, so $n_{0}\leq\tilde{N}_{E}(v,D)$. By construction, the simplicial coefficients of f(C^)f($\widehat${C}) are bounded in absolute value by a constant multiple of the simplicial length $L$ of $\widehat{C}$, and hence by
$\max_{i}|a_{1}^{i}|\leq C_{L}(v,D)\,\operatorname{mass}_{1}(C),$
for some explicit constant $C_{L}(v,D)$.
[…]
@@
$\partial E_{2}=\widehat{C},\ \operatorname{mass}_{2}(E_{2})\leq K_{9}(v,D)\,\sum_{i}|a_{2}^{i}|\leq K_{9}(v,D)\,B_{1}(\tilde{N}_{E}(v,D))\,C_{L}(v,D)\,\operatorname{mass}_{1}(C).$
@@
$\operatorname{mass}_{2}(E)\leq\bigl(C_{2}+C_{\Gamma}C_{1}+K_{9}B_{1}(\tilde{N}_{E})C_{L}\bigr)\operatorname{mass}_{1}(C)\;+\;C_{3}.$
@@
$f_{1}^{\mathrm{Ein}}(v,D):=C_{2}+C_{\Gamma}C_{1}+K_{9}\,B_{1}(\tilde{N}_{E})\,C_{L},$
@@
By construction, all constants involved depend only on the harmonic-radius scale $r_{0}(v,D)$, on the Sobolev constant $C_{S}(v,D)$, on upper bounds for the neck groups $|\Gamma^{i}_{j}|$, and on the bubble–tree complexity $\tilde{N}_{E}(v,D)$, $N_{E}(v,D)$, $k_{E}(v,D)$. ∎

### v2

@@ 2.4. Bubble–tree decomposition
There exist constants $r_{0}(v,D)>0$, $C_{\Gamma}(v,D)>0$ and integers $N_{E}(v,D)$, $k_{E}(v,D)$ together with a decomposition
@@
    (Bodies) If $x\in\mathcal{B}^{i}_{j}$ then
    $r_{h}(x)\geq r_{0}(v,D)\,\operatorname{diam}(\mathcal{B}^{i}_{j}).$
@@
    (Necks) Each $\mathcal{N}^{i}_{j}$ is diffeomorphic to an annulus in $\mathbb{R}^{4}/\Gamma^{i}_{j}$, with $|\Gamma^{i}_{j}|\leq C_{\Gamma}(v,D)$.
@@
This is the specialization of Cheeger–Naber’s bubble–tree construction [8, Theorem 8.63] to the Einstein class. The displayed decomposition above is their formula (8.64). We denote the adjacency relations between bodies and necks as a rooted tree with root $\mathcal{B}^{1}$ and denote by $\mathcal{T}^{i}_{j}$ the subtree with root $\mathcal{B}^{i}_{j}$.
@@ 3.1. Homological filling functions and stationary varifolds
Thus, in order to bound $A_{\min}(M)$ from above, it suffices to obtain a affine upper bound for $\operatorname{FH}_{1}(l)$ in terms of the geometric data of $(M,g)$. The rest of this section is devoted to establishing such a bound in the Einstein setting.
@@ 3.2. A homological filling inequality for Einstein manifolds
A general $1$-cycle $C$ in $M$ will first be homologically replaced by a sum of cycles supported in single bodies, then approximated by a cycle in $\Gamma$, then straightened via the standard graph-to-nerve procedure of [24, Sec. 4] to a simplicial cycle in $\mathfrak{N}$ and filled combinatorially using Lemma 3.7. Each $2$-simplex in $\mathfrak{N}$ corresponds to a geodesic triangle in $\Gamma$, which can be filled by a $2$-chain of uniformly bounded area. This yields the desired affine filling inequality.
@@
where $r_{h}$ is the harmonic radius and $r_{0}(v,D)$ is the constant from Theorem 2.6. Set
@@
The number of sets in $\mathcal{O}$ is bounded by $|\mathcal{O}|\leq\tilde{N}_{E}(v,D)$ for some $\tilde{N}_{E}(v,D)$ depending only on the bubble–tree complexity $N_{E}(v,D)$, $k_{E}(v,D)$ and the covering constants $N_{1}(v,D)$, $N_{2}(v,D)$; see also [24, Sec. 4] for the general Ricci-bounded case.
@@
for some constant $C(v,D)$ depending only on the geometric data. This construction is identical to that in [24, Sec. 4], with the only difference that the scales $R^{i}_{j}$ and $\rho$ are now tied to the harmonic-radius scale $r_{0}(v,D)$ furnished by Theorem 2.6.
@@ 3.3. Local fillings
where $h_{E}(v,D)$ depends only on the uniform bound $C_{\Gamma}(v,D)$ from Theorem 2.6.
@@
The argument is exactly as in [24, Sec. 4.2]. Via $\Phi^{-1}$ we view $C$ as a $1$-cycle in the annulus $A_{\rho,2\rho}(0)\subset\mathbb{R}^{4}/\Gamma$. Radial projection onto a sphere $S_{\lambda}/\Gamma$ with $\lambda\in(\rho,2\rho)$ sweeps out a $2$-chain of area $\lesssim\rho\,\operatorname{mass}_{1}(C)$. The projected cycle is a $1$-cycle in $S^{3}/\Gamma$, whose first homology is finite with generators represented by loops of length $\leq 2\pi\lambda$. Replacing the projected cycle by a representative of its homology class supported in a bounded number of such generators yields $C^{\prime}$ with $\operatorname{mass}_{1}(C^{\prime})\leq h_{E}(v,D)\,\lambda$. Pushing forward by $\Phi$ and controlling the distortion inside the annulus completes the proof. ∎
@@ 3.4. Approximation by $\Gamma$ and improved combinatorial filling
The proof is identical to [24, Lemma 4.5]. One partitions $C$ into short arcs contained in the balls of the cover $\mathcal{O}(\mathcal{B})$, with length $\leq R^{i}_{j}$. Inside each ball one joins the endpoints of an arc to the corresponding center along edges of $\Gamma$. The difference between the original arc and the resulting path in $\Gamma$ lies in a harmonic ball and hence can be filled using Lemma 3.3. Summing over all arcs gives the desired estimates. ∎
The idea of controlling the combinatorial filling via integer solutions of linear systems was introduced in [20]. We show that one can obtain an upper bound for the number of simplices needed to fill a simplicial cycle which depends linearly on the size of the cycle. Note that this is independent of the dimension of the simplicial complex.
@@
As in [24, Lemma 4.6]. If all three vertices lie inside a single body, the triangle lies in a harmonic ball $B_{r_{h}/2}$ and can be filled using Lemma 3.3. If they lie inside a single neck, the triangle lies in the union of at most two trapezoids contained in $\bar{T}_{i,j,\ell}$, and can be filled by a cone in the annulus, with area $\lesssim\rho^{2}$. Mixed configurations are reduced to these two cases. All constants can be controlled in terms of $r_{0}(v,D)$ and $D$. ∎
@@ 3.5. Decomposition and the Einstein filling inequality
As in [24, Sec. 4.3], one cuts $C$ whenever it crosses a boundary component $\partial\mathcal{B}\cap\mathcal{N}$ and reassembles the resulting arcs into cycles supported in individual bodies or necks. The number of such regions is bounded by the bubble–tree complexity $N_{E}(v,D)\,k_{E}(v,D)$, and the total length increases by at most a uniform multiplicative factor $C_{0}(v,D)$, coming from bounded overlap of the decomposition and the uniform control on the diameter of bodies and necks. ∎
@@
$\partial E_{0}=C-C^{\mathrm{body}},\quad\operatorname{mass}_{2}(E_{0})\leq C_{2}(v,D)\,\operatorname{mass}_{1}(C)+C_{3}(v,D),$
@@
The remaining step is the chain-level passage from the graph $\Gamma$ to the nerve $\mathfrak{N}$. Instead of using the partition-of-unity map directly on $\widehat{C}$, we invoke the graph-to-nerve straightening procedure from [24, Sec. 4]. Applied to the graph cycle $\widehat{C}$, it yields a simplicial $1$-boundary
$c_{1}=\sum_{i}a_{1}^{i}\,\sigma_{i}^{1}$
in $\mathfrak{N}$ and a $2$-chain $E_{\mathrm{st}}$ in $M$ such that
$\partial E_{\mathrm{st}}=\widehat{C}-\gamma(c_{1}),$
where $\gamma(c_{1})$ denotes the edge cycle in $\Gamma$ corresponding to $c_{1}$, and
$\sum_{i}|a_{1}^{i}|\leq C_{L}(v,D)\,\operatorname{mass}_{1}(C),\qquad\operatorname{mass}_{2}(E_{\mathrm{st}})\leq C_{L}(v,D)\,\operatorname{mass}_{1}(C).$
In particular,
$\max_{i}|a_{1}^{i}|\leq C_{L}(v,D)\,\operatorname{mass}_{1}(C).$
Let $n_{0}$ be the number of vertices of $\mathfrak{N}$, so $n_{0}\leq\tilde{N}_{E}(v,D)$. Since $c_{1}$ is a simplicial boundary in $\mathfrak{N}$, Lemma 3.7 with $k=1$ yields a simplicial $2$-chain $C_{2}=\sum a_{2}^{i}\sigma_{i}^{2}$ with $\partial C_{2}=c_{1}$ and
@@
$\partial E_{2}=\gamma(c_{1}),\ \operatorname{mass}_{2}(E_{2})\leq K_{9}(v,D)\,\sum_{i}|a_{2}^{i}|\leq K_{9}(v,D)\,B_{1}(\tilde{N}_{E}(v,D))\,C_{L}(v,D)\,\operatorname{mass}_{1}(C).$
@@
$\operatorname{mass}_{2}(E)\leq\bigl(C_{2}+C_{\Gamma}C_{1}+C_{L}+K_{9}B_{1}(\tilde{N}_{E})C_{L}\bigr)\operatorname{mass}_{1}(C)\;+\;C_{3}.$
@@
$f_{1}^{\mathrm{Ein}}(v,D):=C_{2}+C_{\Gamma}C_{1}+C_{L}+K_{9}\,B_{1}(\tilde{N}_{E})\,C_{L},$
@@
[…]